OpenStudy (anonymous):

find all roots of x^6-13^4+39x^2-27=0

OpenStudy (anonymous):

x^6-13^4+39x^2-27=0 ^ x there?

OpenStudy (anonymous):

factor again if you missing the x there (x-1)*(x+3)*(x-3)*(x+1)*(x^2-3)

OpenStudy (anonymous):

[x = 1], [x = -3], [x = 3], [x = -1], [x = sqrt(3)], [x = -sqrt(3)]

OpenStudy (anonymous):

no their is an x right after the 13^4

OpenStudy (anonymous):

you missing the x Jean :)

OpenStudy (anonymous):

yes I know, it is x^6-13^4x+39x^2-27=0

OpenStudy (anonymous):

(x-1)*(x+3)*(x-3)*(x+1)*(x^2-3) =0

OpenStudy (anonymous):

So would it be 1,-3,3,-1.

OpenStudy (anonymous):

yes and sqrt(3) -sqrt(3)

OpenStudy (anonymous):

YOU FORGET (x^2-3) =0 add 3 x^2=3 take sqrt x= plus or minute sqrt(3) -sqrt(3) sqrt(3)

OpenStudy (anonymous):

plus or minus*

OpenStudy (anonymous):

Um so the answer is 1,sqrt(3), sqrt(3),-1, and x=plus or minus 3. I am a little confused.

OpenStudy (anonymous):

1,-3,3,-1. are roots but also -sqrt(3) and sqrt(3)

OpenStudy (anonymous):

okayyy. I get it now.

OpenStudy (anonymous):

Good job Marge :D